CH/Agen/92/06

Agenda

Charissa Group Meeting

To be held on: 25 November 1992, 11.30 am To be held at: York

Distribution: WNC,NMC,RAC,BRF,GJG,SJAH,RAH,JSL,ACM, VP,WDMR,GT,RPW,DLW

1.  Minutes of Meeting of 7 October 1992
2.  Matters Arising/ Actions
3.  Update on NSC/SERC matters  (DLW)
4.  Progress on silicon detectors  (BRF)
5.  Progress on scintillators  (NMC)
6.  Progress on chamber & detector assemblies  (JSL)
7.  Progress on gas supply  (RAC)
8.  Progress on electronics & data acquisition  (RAH)
9.  Developments on hybrid detectors  (RAC)
10.  Developments on event manager & GEC emulator  (VP)
11.  Documentation of Experimental Setup  (BRF)
12.  Completion of Programme at NSF  (BRF)
13.  Responsibilities for ANU Preparation  (BRF)
14.  Student Training  (DLW)
15.  Future of DL Detector Laboratory
16.  Disposal of NSF radioactive sources
17.  Disposal of NSF “minor” equipment
18.  Any other business
19.  Date of next meeting

W.N. Catford (Secretary) 18 October 1992 Item 11: Documentation

From BRF: At the recent run (^{24}Mg+ ^{24}Mg) it became clear that we need to collect together

and update the documentation on Charissa experiments and make it available to all. We have had information available in the past, but it has been prepared on a rather ad hoc basis and not everyone has received copies. With the large number of new people joining the collaboration, and the addition of new equipment (gas detectors and strip detectors) there is a need to prepare an up-to-date and comprehensive set of notes covering all aspects of the experiments. This will be particularly useful for those out at the ANU during the setup phase.

I would like to see a single manual containing sections on all areas of the experiment. A copy would be available for every member of the group.

A suggested contents list is:

• Chamber: beamline, collimation,detector mounts, cables, chamber control

• Electronics: setups for strip-strip runs and hybrid-hybrid runs

• Data Acquisition: hardware setup and software control

• Running: checklist of proceedures (stages to go through pre-beam, during calibration runs and during reaction runs)

ACTION: Discuss this. Agree contents. Delegate people to produce drafts. Set timetable. (WDMR has already agreed to do 2)

Item 12: Completion of Programme at NSF

From BRF: It now appears likely that we will be able to fit in two more runs on the NSF

before it closes (with a few extra shifts added to each run for detector tests). At present we have four approved experiments:

Seven alpha chain state High resolution run Angular distribution measurement C—C* decay search

and we need to decide which to plan for. One of the main considerations is to ensure everyone (esp. students) has data to work on between March and September (when we run at the ANU).

This year we have had new students starting at York (Steve Chappel), Surrey (Steve Singer) and Birmingham (Neil Curtis) as well as a new RA at Birmingham (Roger Ward). Steve Chappel will have the Orsay experiment to plan for and analyse, and Steve Singer has started on the analysis of the ^{12}C+^{24}Mg experiment.

I would suggest, therefore, that we run the high resolution experiment for Neil Curtis and the C—C* search for Roger Ward.

The alpha chain experiment and the angular distribution measurement then form the basis of our first running period at the ANU (plus any other experiments which arise between now and then).

There is likely to be a recommendation to the scheduling committee that Charissa be given priority for any additional NSF time (for example if an experiment fails or if the machine performs better than anticipated). We need to be ready to step in with another run and should decide which to go for.

ACTION: Discuss these suggestions and agree action.

Item 13: Responsibility for ANU Preparations

From BRF:

A number of things need to be done in preparation for our running at the ANU. The jobs at the moment are as listed below, but more will undoubtedly arise in the following months.

1. New detector mounts and collimation for chamber

2. New chamber feedthroughs

3. New cable runs from chamber to control room

4. Electronics and DA shipped out and installed

5. Beamline acquired and shipped out.

In order to ensure these are carried out it would seem sensible for one group to take responsibility for a particular job. I would offer the following suggestions:

1. York (machine shop personel)

2. York (ditto)

3. Surrey (most familiarity with layout)

4. Birmingham/Oxford (Experience in shipping and setting up)

5. Birmingham (ditto)

ACTION: Discuss these ideas. Identify any other jobs. Agree responsibilities. Agree completion dates.